OpenStudy (anonymous):

find solution in R in the interval )-π,π). sin3x= -sin2x

ganeshie8 (ganeshie8):

use "sinc + sind" formula

ganeshie8 (ganeshie8):

SinC + SinD = 2Sin[(C+D)/2]Cos[(C-D)/2]

ganeshie8 (ganeshie8):

sin3x= -sin2x sin3x + sin2x = 0

ganeshie8 (ganeshie8):

apply the formula now

OpenStudy (anonymous):

but than what am goona do later

ganeshie8 (ganeshie8):

yeah lets see... first, wat do u get after applying the formula ?

ganeshie8 (ganeshie8):

sin3x= -sin2x sin3x + sin2x = 0 2sin(5x/2) cos(x/2) = 0

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

wait am trying to get

ganeshie8 (ganeshie8):

take ur time :)

OpenStudy (anonymous):

2sin(5x/2) cos(x/2)=0!!!!

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

can u explain more

ganeshie8 (ganeshie8):

sure :) next :- 2sin(5x/2) cos(x/2)=0 sin(5x/2) = 0 OR cos(x/2) = 0

ganeshie8 (ganeshie8):

you need to solve both equations okay ?

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

2sin(5x/2) cos(x/2)=0 sin(5x/2) = 0 OR cos(x/2) = 0 \(\large \frac{5x}{2} = \sin^{-1}0\) OR \(\large \frac{x}{2} = \cos^{-1}0\)

ganeshie8 (ganeshie8):

sin = 0 for what angels ?

OpenStudy (anonymous):

kp

ganeshie8 (ganeshie8):

whut.. ?

OpenStudy (anonymous):

x=kπ

ganeshie8 (ganeshie8):

yes, sin(x) is 0, whenever x = kpi

ganeshie8 (ganeshie8):

2sin(5x/2) cos(x/2)=0 sin(5x/2) = 0 OR cos(x/2) = 0 \(\large \frac{5x}{2} = \sin^{-1}0\) OR \(\large \frac{x}{2} = \cos^{-1}0\) \(\large \frac{5x}{2} = k \pi\) OR \(\large \frac{x}{2} = ?????\)

ganeshie8 (ganeshie8):

wat about cos ? cos = 0 for what angles ?

OpenStudy (anonymous):

x=,π/2+k,π right!!

ganeshie8 (ganeshie8):

yup, cos(x) = 0, when x = pi/2 + kpi

ganeshie8 (ganeshie8):

2sin(5x/2) cos(x/2)=0 sin(5x/2) = 0 OR cos(x/2) = 0 \(\large \frac{5x}{2} = \sin^{-1}0\) OR \(\large \frac{x}{2} = \cos^{-1}0\) \(\large \frac{5x}{2} = k \pi\) OR \(\large \frac{x}{2} = \frac{\pi}{2} + k \pi\)

ganeshie8 (ganeshie8):

so, the general solutions are :- \(\large \frac{5x}{2} = k \pi\) OR \(\large \frac{x}{2} = \frac{\pi}{2} + k \pi\)

OpenStudy (anonymous):

thank uu :)

ganeshie8 (ganeshie8):

we're not done yet, we need to find "particular solutions" in the interval -pi and pi

ganeshie8 (ganeshie8):

so, the general solutions are :- \(\large \frac{5x}{2} = k \pi\) OR \(\large \frac{x}{2} = \frac{\pi}{2} + k \pi\) \(\large x = \frac{2}{5}k \pi\) OR \(\large x = \pi + 2k \pi\)

ganeshie8 (ganeshie8):

above is the general solution, to get particular solutions in interval (-pi, pi) :- put k = -2, -1, 0, 1, 2... and pick the x values that are between -pi and pi

OpenStudy (anonymous):

yess i know how

ganeshie8 (ganeshie8):

you should get below solutions :- \(\large \frac{-4\pi}{5}, \frac{-2\pi}{5}, 0, \frac{2\pi}{5} , \frac{4\pi}{5}\)

ganeshie8 (ganeshie8):

incase if u want to check ur answer... :)

OpenStudy (anonymous):

:) (y)

ganeshie8 (ganeshie8):

u may check wid wolfram also : http://www.wolframalpha.com/input/?i=sin%283x%29%3D+-sin%282x%29%2C+-pi%3Cx%3Cpi

OpenStudy (anonymous):

why nott if i want help can ask u again????

ganeshie8 (ganeshie8):

sure :))